I need help with the algebra if the problem. I saw solutions showing the correct final answer but I do not get the final answer they get. Please show the algebra of the last third of the problem. So I understand what do I do wrong.

Answer: 0.587 rad/s^21. The 50-kg uniform crate rests on the platform for which the coefficient of static friction is us = 0.5. If the supporting
Answer #1

Mass of crate= 50 kg. Coefficient of static flichen, = 0.5 . Tangential acceleration, 9+ = 910. Normal acceleration, an= 99 w0.5m. man. LW 1 sm 331.85-100 = 173.22 +50d. 231.85 = 223-22 2 = 1,0387 rad /s2 Case (2) when Gate is on the verge of sipping